BMW is taking on new project and came up with the following projected cash flows. Compute a) Profitability Index b) NPV c) IRR for the project. Use a discount rate of 10%. Discuss the results by explaining what each of the methods mean and your decision on going forward with the project.

Year

Cash flow ($)

0

(850,000)

1

350,000

2

350,000

3

250,000

4

200,000

5

200,000
Answer #1

First let us find out the present values of cash outflows and cash inflows:

Now that we have present value of cash outflows and inflows, let us find the (a) Profitability Index (PI)

Profitability~Index= \frac{Present~Value~of~Total~Cash~inflows}{Present~Value~Total~Cash~outflows}

= \frac{1,056,053.67}{850,000}

= 1.24

b) Net Present Value (NPV):

Net~Present~Value= Present~Value~of~Total~Cash ~inflows - Present~Value~of~Total~Cash~outflows

= 1,056,053.67 - 850,000

= 206,053.67

c) Internal Rate of Return (IRR):

We can use trial and error method here, we have to select 2 discount rates where, when discounted one should show greater than cash outflow and other should show less than cash inflow. We cannot use 10% here because it is too far from 850,000 which can give you errors.

Here we have taken 20% and 22%

Cash flows Cash flows Discount Factor Cash flows Discount Factor @ 20% Discounted at 20% Discounted at 22% Year @ 22% 0.83333

IRR=20 + \left [ \frac{856224.28-850000}{856224.28-823993.48} \right ]*(22-20)

= 20.39

Answer Meaning It is used for ranking the projects. When profitability index is above 1, it signals to accept the project. In